# INACIA: Integrating Large Language Models in Brazilian Audit Courts: Opportunities and Challenges

### 19 Jan 2024 v2

@@ 3.3.2. Extraction of Allegations and Requests
Figure 3. Prompt used to extract the allegations and requests.
@@
Similar to the basic information extraction task, we also implemented a combination of prompt engineering plus RAG to extract the relevant information (see Figure 3). We provide the model with an example of the allegations and requests to be extracted. The model extracts the allegations and requests from the case’s representação. We used OpenAI’s GPT-4 model with a context length of 32,000 tokens. The main advantage of this model is that it does not require fine-tuning or any additional training. The prompt contains instructions, an example of the allegations and requests to be extracted, a set of rules the LLM must follow, and the relevant case documents.
@@ 3.4. Admissibility Examination
The next step in the INACIA project is to conduct an admissibility examination (“Admissibility exam” blue box in Figure 1). The admissibility examination is a preliminary examination conducted by auditors to determine whether the claims meet the minimum legal criteria to proceed. The TCU defines five criteria that must be met for a claim to be admissible:
@@
As shown in Figure 4, we used an approach similar to the one used for basic information extraction. We applied an LLM as an intelligent agent capable of reasoning and interpreting legal documents. However, instead of feeding the LLM the first two pages of the case’s representação, we allow the model to search for information in case documents using a search engine – including four external databases: (1) TCU’s jurisprudence database; (2) TCU’s statutes and Federal law; (3) TCU’s internal codes and other documents; and (4) the case’s documents.
@@
In this task, the agent can compose queries to search for information in case documents and external databases. For example, the agent may want to compose the query quem tem legitimidade para apresentar uma representação ao TCU? (i.e., “who has legitimacy to file a claim to the TCU?”) to search for information about who can file a claim in external databases. If that’s the case, the agent should not be limited to pre-defined queries and keyword-based queries. To facilitate this, we employ NeuralSearchX ((Almeida et al., 2022)), a search engine that integrates the capabilities of the BM25 algorithm and a neural reranker (Nogueira et al., 2020), fine-tuned for Brazilian Portuguese. NeuralSearchX first retrieves documents containing the words from the query and then utilizes its re-ranking mechanism to prioritize the most relevant documents.
In addition to searching, we used an approach based on the Visconde system (Pereira et al., 2023) to reason through the information retrieved from case documents and external databases. Visconde is a multi-document question-answering system integrating information retrieval and LLMs to answer questions. The reasoning step is based on the Chain-of-Thought approach (Kojima et al., 2022) and produces a paragraph explaining how the documents support the model’s answer.
@@ 3.5. Periculum in mora Analysis
The analysis, however, is relatively straightforward. We carry out a simple search for exact matches of minuta de contrato (draft contract) in all case documents after a case has been assigned to a TCU staff, except when the case contains fewer than five documents – in which case all documents are considered. This search is essential to flag documents mentioning the draft contract so they are not included in subsequent steps. Two key rules are used to standardize the decision on whether there is Periculum in mora: (1) the existence of an active or signed contract – there is no danger in delay since TCU cannot prevent a contract from coming into force; (2) and (2) the existence of events that prove delay, cancellation, or impediment in contracting – there is no danger in delay because the execution of a contract has been halted. The verification of these rules occurs step-wise. We first verify whether there is any indication that a goods or service provision contract has been signed or is active. The verification is done directly from a keyword search and indirectly by asking the LLM to reason through the documents and verify the existence and status of procurement contracts. The second step is verifying any event delaying, canceling, or suspending a contract from entering into force. Here, we do not perform a direct keyword search for such information; we send documents and prompts straight to the LLM. The outcome of these searches provides an answer regarding the existence of an active contract and any delays or impediments. Finally, the findings are synthesized into a dictionary containing information about the presence of an active contract or delays, along with the document IDs from each search and a suggested decision on whether there is Periculum in mora, which can be either “rejected” or “accepted”.
@@ 3.6. Fumus boni iuris Analysis
For the Fumus boni iuris analysis, we used an approach similar to the one used for the admissibility examination. We used the LLM as an intelligent agent capable of reasoning and interpreting legal documents. However, instead of feeding the LLM with one of the admissibility criteria, we feed the LLM with the allegations extracted in the pipeline task “allegations and requests”. The LLM then searches for information in case documents and external databases to determine whether there is Fumus boni iuris.
The process for the Fumus boni iuris analysis is divided into two steps. In the first step, we feed the agent the allegations extracted in the previous project stage and ask him to classify each allegation as grounded in law, not grounded in law, or inconclusive. We use an approach based on the ReAct framework (Yao et al., 2023). We use the reasoning and acting components of the ReAct framework to instruct the model to classify the allegations. The reasoning component is used to reason through the allegations and plan the actions to be taken. The acting component executes the actions planned by the reasoning component, which includes searching the case documents and the external databases. One of the external databases used in this project stage is TCU’s selection of statutes and Federal law. The model searches for information about the laws invoked by the allegations in this database. The agent also searches for information about the allegations in the TCU’s jurisprudence database. The model then classifies the allegations as grounded in law, not grounded in law, or inconclusive.
@@ 3.7. Recommendations
The INACIA pipeline concludes by generating a paragraph that, in this context, is called recommendations. The recommendations are a set of directions for the decision-making authority to address the case. As the text is the auditor’s responsibility, INACIA only suggests directions for the auditor to consider. This pipeline task is relatively straightforward and recasts the other outputs from each task to construct a cohesive and coherent instruction document to be sent over to the panel of judges who will issue the final ruling on a case. To write case recommendations, INACIA uses the following components:
@@ 4. Experiments
This section describes the experiments we conducted to evaluate the INACIA pipeline. Although INACIA has a variety of project tasks, we only evaluated the recommendations task, which represents the system’s final and most critical output. We first describe the dataset we used in the evaluation and then describe the evaluation techniques we used. Finally, we describe the results of the experiments.
@@ 4.1. Dataset Creation and Curation
A validation dataset was created to assess the quality of the analyses performed by INACIA. This dataset consists of a fully annotated table with the expected results for each pipeline step. In this study, we only worked with closed cases. This means the document of interest for evaluation (the initial instruction document) was already available for all cases. This document is a searchable PDF written by TCU staff with expertise on the case subject. It is written in natural language and divided into five main sections: basic information, claims, and requests, admissibility examination, precautionary measure analysis (periculum in mora and fumus boni iuris, and recommendations. We used instructions from closed cases to build the validation dataset.
We extracted various elements from the instruction text to build the validation database. Each element followed its extraction strategy, as the information had different data structures. For basic information, GPT-4 was used to extract the data, such as names of parties, the amount involved, contract duration, and case summary, among others. In total, 26 basic variables were extracted. Before using GPT-4 in all cases, a random sample of 30 cases was manually analyzed to check if the data was being extracted with quality. This analysis resulted in 100% accuracy for the main information of interest.
For the claims and requests piece, the text was directly extracted from the instruction document, using regular expressions to isolate the section of interest. As the instruction documents considered are standardized, this extraction presented no errors. The same strategy was applied to the recommendations.
In the case of the admissibility examination, there are five criteria of interest, as described in Section 3.4. For each criterion, the labels “Yes (met)”, “No (not met)”, “Partial (partially met)”, or “Not applicable” were extracted using regular expressions. This information is systematically annotated in the instructions; hence, no extraction errors or manual corrections were needed.
The same technique of the admissibility examination was applied to the analysis of granting precautionary measures. There are three main criteria of interest: periculum in mora (danger in delay), reverse periculum in mora, and fumus boni iuris (legal plausibility). In all these cases, the possible outcomes of the analysis are “Dismissed (rejected)”, “Configured (accepted)”, “Inconclusive”, or “Not analyzed” (for example, when the case is not known). These data were extracted using regular expressions and did not present extraction errors.
Finally, a validation table was produced for all the cases of interest, containing basic information, claims, and requests, admissibility examination, analysis for granting precautionary measures and forwarding proposals. The final validation database effectively considered was a sample of 122 cases from all the over 1000 cases analyzed. The selection of cases that would be part of the final validation table was based on both some additional scope filters, focusing on cases with simpler analysis, and a random sample of the cases resulting from the filters.
@@
Although we collect and validate a wide range of information, including basic case details, claims, admissibility assessments, and so on, not all results carry equal significance. We emphasize the evaluation of the recommendations, as it is the last and most important step of the instruction process.
@@ 4.2. Evaluation
For evaluating the INACIA main output, we used the validation dataset described in the previous section. The evaluation aims to compare the outcomes produced by the INACIA pipeline with the expected results previously produced by human auditors. This comparison could be done using term-matching metrics for these tasks, such as BLEU, Rouge, etc. However, these metrics are unsuitable for these tasks because the text generated by INACIA may not be the same as the text in the reference form. Instead, we used the LLM to evaluate if the candidate text is consistent with the reference text. Instead of asking the model to classify the candidate text as consistent or inconsistent with the reference text, we employ a more comprehensive approach.
@@
Evaluating text generated by LLMs is a challenging task. While metrics such as BLEU and ROUGE are commonly used to evaluate text generation tasks, they have shown low correlation with human judgment (Fabbri et al., 2021). Approaches like GPTScore (Fu et al., 2023), and G-eval (Liu et al., 2023) have been proposed to address this issue by using LLMs to evaluate the generated text. They have been shown to correlate more with human judgment than traditional metrics when evaluating text summarization. However, these approaches are unsuitable for evaluating the text generated by the INACIA pipeline because the generated text is not a summary of a document. Instead, the generated text is the auditor’s recommendations.
@@
For example, the objective of evaluating the recommendation generation is to assess whether the recommendations produced by INACIA are consistent with those produced by a human auditor. Intuitively, we want to evaluate if the candidate text (INACIA) contains the same elements as the reference text (auditor). For this, we prompted OpenAI’s GPT-4 with the reference recommendations from the validation dataset, asking it to compose a checklist of the elements that should be present in the recommendations. We then prompted the model with the candidate recommendations and asked it to evaluate if they contained the checklist elements. Then, we can count the number of checklist elements present in the candidate recommendations and use this number to evaluate the quality of the extracted information. For example, if the checklist contains ten elements and the candidate recommendations contain eight elements, we can say that the candidate recommendations are 80% consistent with the reference recommendations.
@@
We also used the candidate document to prompt the model to compose a checklist of the elements that should be present in the candidate document. We then prompted the model with the reference document and asked it to evaluate if it contained the checklist elements. Then, we can count the number of checklist elements present in the reference document and use this number to evaluate the quality of the extracted information. This way, we have two scores: the precision, which indicates how consistent the candidate document is with the reference document, and the recall, which indicates how consistent the reference document is with the candidate document. We then use the F1 score to evaluate the quality of the extracted information. The F1 score is the harmonic mean of the precision and recall. We used the same approach to evaluate the recommendation generation.
@@
This approach has shown to have a good performance in assessing the quality of the extracted text in the TCU context. To evaluate it, we asked an auditor to analyze five examples in which the score is too low (less than 0.3) and too high (more than 0.7). The auditor concluded that the examples with low scores were not consistent with the reference text, and those with high scores were consistent with the reference text. This shows that the approach is suitable for evaluating the quality of the extracted text in the TCU context. Yet, we evaluated this approach on assessing the consistency of the automatically generated summary so we could compare it with GPTScore and G-eval. We used the SummEval Dataset (Fabbri et al., 2021) as a benchmark. The results shown in Table 1 demonstrate that our approach correlates more with human judgment than GPTScore and G-eval on assessing how consistent a summary is concerning its original text. It also has a higher correlation with human judgment, considering the overall quality of the summary. This shows that our approach is suitable for evaluating the quality of the automatically generated text in the TCU context.

### 26 Feb 2024 v3

@@ 4.3.2. Extraction of Allegations and Requests
Figure 4. Prompt used to extract the allegations and requests in a representação.
In Figure 4, we present the prompt used at this pipeline step. It has four main characteristics: (1) the instruction (identified by the title “## Instruction”), which tells the model its role and what we want it to do; (2) the example (identified by the title “## Example”), which teach the model the expected structure and language of the text; (3) the rules (identified by the title “## Rules”), which are the constraints that the model must follow to generate the extract the allegations and requests; and (4) the input, i.e., the case’s main document, from which the model will extract the allegations and requests.
For this step, we used OpenAI’s GPT-4 model with a context length of 32,000 tokens. The main advantage of this model is that it does not require fine-tuning or additional training. Besides, it was trained to generate human-like text, which is essential for the INACIA project. The high context length is also important because the documents can be long and complex, and the model needs to understand the context to generate the allegations and requests correctly.
@@ 4.4. Admissibility Examination
The next step in the INACIA project is to conduct an admissibility examination (“Admissibility exam” blue box in Figure 1). The admissibility examination is a preliminary examination conducted by auditors to determine whether the claims meet the minimum legal criteria to proceed (i.e., whether TCU should hear the case). TCU defines five criteria that must be met for a claim to be admissible:
@@
As shown in Figure 5, we used an approach similar to the one used for basic information extraction. We applied an LLM as an intelligent agent capable of reasoning and interpreting legal documents. However, instead of feeding the LLM with the first two pages of the case’s representação, we allow the model to search for information in case documents using a search engine – including four external databases: (1) TCU’s jurisprudence database; (2) TCU’s statutes and Federal law; (3) TCU’s internal codes and other documents; and (4) the case’s documents.
@@
In this task, the agent can compose queries to search for information in case documents and external databases. For example, the agent may want to compose the query quem tem legitimidade para apresentar uma representação ao TCU? (i.e., who has legitimacy to file a claim to the TCU?) to search for information on who can file a claim in external databases. If that’s the case, the agent should not be limited to pre-defined queries and keyword-based queries. To facilitate this, we employ NeuralSearchX (Almeida et al., 2022), a search engine that integrates the capabilities of the BM25 algorithm and a neural reranker (Nogueira et al., 2020), fine-tuned for Brazilian Portuguese. NeuralSearchX first retrieves documents containing the words from the query and then utilizes its re-ranking mechanism to prioritize the most relevant documents.
In addition to searching, we used an approach based on the Visconde system (Pereira et al., 2023) to reason through the information retrieved from case documents and external databases. Visconde is a multi-document question-answering system integrating information retrieval algorithms and LLMs to answer questions. The reasoning step is based on the Chain-of-Thought approach (Kojima et al., 2022) and produces a paragraph explaining how the documents support the model’s answer.
@@ 4.5. Periculum in mora Analysis
For this analysis, we carry out a search for exact matches of minuta de contrato (draft contract) in all case documents after a case has been assigned to a TCU staff, except when the case contains fewer than five documents – in which case all documents are considered. This search is essential to flag documents mentioning the draft contract so they are not included in subsequent steps. Two key rules are used to standardize the decision on whether there is Periculum in mora: (1) the existence of an active or signed contract – there is no danger in delay since TCU cannot prevent a contract from coming into force; and (2) the existence of events that prove delay, cancellation, or impediment in contracting – there is no danger in delay because the execution of a contract has been halted. The verification of these rules occurs step-wise. We first verify whether there is any indication that a goods or service provision contract has been signed or is active. The verification is done directly from a keyword search and indirectly by asking the LLM to reason through the documents and verify the existence and status of procurement contracts. The second step is verifying any event delaying, canceling, or suspending a contract from entering into force. Here, we do not perform a direct keyword search for such information; we send documents and prompts straight to the LLM. The outcome of these searches provides an answer regarding the existence of an active contract and any delays or impediments. Finally, the findings are synthesized into a dictionary containing information about the presence of an active contract or delays, along with the document IDs from each search and a suggested decision on whether there is Periculum in mora, which can be either “rejected” or “accepted”.
@@ 4.6. Fumus boni iuris Analysis
For the Fumus boni iuris analysis, we used an approach similar to the one used for the admissibility examination. We used the LLM as an intelligent agent capable of reasoning and interpreting legal documents. However, instead of feeding the LLM with one of the admissibility criteria, we feed the LLM with the allegations extracted in the pipeline task “allegations and requests” (cf. Section 4.3.2). The LLM then searches for information in case documents and external databases to determine whether there is Fumus boni iuris.
The process for the Fumus boni iuris analysis is divided into two steps. In the first step, we feed the agent the allegations extracted in the previous project stage and ask him to classify each allegation as grounded in law, not grounded in law, or inconclusive. We use the reasoning and acting components of the ReAct framework (Yao et al., 2023) to instruct the model to classify the allegations. The reasoning component is used to reason through the allegations and plan the actions to be taken. The acting component executes the actions planned by the reasoning component, which includes searching the case documents and the external databases. One of the external databases used in this project stage is TCU’s selection of statutes and Federal law. The model searches for information about the laws invoked by the allegations in this database. The agent also searches for information about the allegations in the TCU’s jurisprudence database. The model then classifies the allegations as grounded in law, not grounded in law, or inconclusive.
The second step consists of producing a decision about the Fumus boni iuris of the case. The decision is based on the classification of the allegations. For this step, we prompt the LLM to classify the allegations in three classes: inconclusivo (inconclusive), afastada (not grounded in law), and configurada (grounded in law). We provide the model with the allegations and the classification of each allegation obtained in the first step. We then ask the model to produce a decision about the Fumus boni iuris of the case. The model uses the classification of the allegations to produce the decision. Before the decision is made, the model is instructed to produce a paragraph explaining the decision. The paragraph should explain the decision and the reasons for the decision. The paragraph is included in the initial instruction document.
@@ 4.7. Recommendations
The INACIA pipeline concludes by generating a paragraph that, in this context, is called recommendations (i.e., the adjudication suggestions). The recommendations are a set of directions for the decision-making authority to address the case. Since the auditor assigned the case bears full responsibility for the adjudication recommendations, INACIA only suggests directions for the auditor to consider but the auditor has the final say on whether they accept the suggestion. This pipeline task is relatively straightforward and recasts the other outputs from each task to construct a cohesive and coherent instruction document to be sent over to the panel of judges who will issue the final ruling on a case. To write adjudication recommendations, INACIA uses the following components:
@@ 5. Experiments
This section describes the experiments we conducted to evaluate the INACIA pipeline. Although INACIA has a variety of project tasks, we only evaluated the recommendations task, which represents the system’s final and most critical output. We first describe the dataset we used in the evaluation and then describe the evaluation techniques we used. Finally, we present the results of the experiments.
@@ 5.1. Dataset Creation and Curation
A validation dataset was created to assess the quality of the analyses performed by INACIA. This dataset consists of fully annotated cases for all information pieces extracted at each step of the pipeline. In total, we annotated 122 cases and 33 variables per case.777 The selection of these cases followed criteria set up by the research team and TCU to better evaluate INACIA given the wide variety of cases processed by the court: we blocked cases whose claims were violations of one procurement law, between 2019 and 2022, and which contained periculum in mora analyses (such that we could recreate the complete analysis needed by TCU). Though the final evaluation is only performed with the recommendations task, we tracked all other variables for intermediate outputs and for adjusting each INACIA module to output the necessary results for the following module. The annotation of these cases consisted primarily of having the language model extract each variable from the initial instruction of every case. Since we only worked with disposed cases, we had the adjudication recommendation document (i.e., the initial instruction) available to us when creating this validation dataset, which served then as the single source of validation information. The split of extracted variables is (cf. Appendix A):
  • 23 variables in the basic information module;
  • 5 variables in the admissibility examination module;
  • 2 variables in the periculum in mora module;
  • 1 variable in the fumus boni iuris module;
  • 1 variable in the allegations and requests module;
  • 1 variable in the recommendations module.
We instructed GPT-4 to extract all the information above from the initial instruction of each case to retrieve the labels INACIA would have to predict based on case filing documents (i.e. the representação and supporting documents). [^8] This process is exactly what TCU auditors do: from a set of documents filed by the plaintiff( i.e., the representação and supporting documents) they draft the initial instruction. To guarantee that the validation dataset was being constructed correctly, we assigned one senior researcher in our team (Ph.D. level) to manually review 30 randomly sampled cases for all variables to then compare the researcher’s data extraction to that of GPT. GPT achieved 100% accuracy on this task, and we confidently proceeded to ask it to code the 122 selected cases.
@@
Although we collect and validate a wide range of information, including basic case details, claims, admissibility assessments, and so on, the results do not carry equal significance. We emphasize the evaluation of the recommendations, as it is the last and most important step of the instruction process.
@@ 5.2. Evaluation
For evaluating the INACIA main output, we used the validation dataset described in the previous section. The evaluation aims to compare the outcomes produced by the INACIA pipeline with the expected results previously produced by human auditors. We only evaluated INACIA on the recommendations task (cf. Section 4.7), which is the consolidation of outputs of all intermediate tasks (listed in Sections 4.3 to 4.6) to write the initial instruction document.
@@
The evaluation of recommendations could be done using term-matching metrics, such as BLEU, ROUGE, etc. However, evaluating free-form text generated by LLMs is a challenging task. While BLEU and ROUGE are commonly used to evaluate text generation tasks, they have shown low correlation with human judgment (Fabbri et al., 2021). Approaches like GPTScore (Fu et al., 2023), and G-eval (Liu et al., 2023) have been proposed to address this issue by using LLMs to evaluate the generated text. They have been shown to correlate more with human judgment than traditional metrics when evaluating text summarization. However, these approaches are also unsuitable for evaluating the text generated by the INACIA pipeline because the generated text is not a summary of a document. Instead, the generated text is similar to what the auditor’s adjudication recommendations were.
@@
The objective of this evaluation thus is to assess whether the recommendations produced by INACIA are consistent with those produced by a human auditor. Intuitively, we want to evaluate if the candidate text (INACIA) contains the same elements as the reference text (human auditor). For this, we used the approach depicted in Figure 6. First, we prompted OpenAI’s GPT-4 with the reference recommendations from the validation dataset, asking it to compose a checklist of the elements that should be present in the recommendations (i.e., the Checklist Generation step). The result is a list of elements that should be present in the recommendations (i.e., the evaluation checklist). Each element is a sentence with a boolean statement (e.g., “The case is admissible”). We then prompted the model with the candidate recommendations alongside the evaluation checklist and asked it to evaluate if the text contained the checklist elements (i.e., the Checklist Evaluation step). Then, we can count the number of checklist elements present in the candidate recommendations and use this number to evaluate the quality of the extracted information. This produces a score that indicates how consistent the candidate recommendations are with the reference recommendations. For example, if the checklist contains ten elements and the candidate recommendations contain eight elements, we can say that the candidate recommendations are 80% consistent with the reference recommendations.
Figure 6. The checklist approach used to evaluate the recommendations.
Figure 7 shows an example of a checklist of elements that should be present in the recommendations of an example case. It was composed by OpenAI’s GPT-4 originally in Brazilian Portuguese and then translated to English for this paper. The checklist contains five elements that should be present in the candidate text. This way, we can evaluate the quality of the recommendations by comparing the checklist elements with the candidate recommendations. It is worth noting that the number of elements is not fixed and can vary from case to case.
  1. (1)
    Verificar se a representação atende aos requisitos de admissibilidade conforme Lei 8.666/1993, art. 113, § 1º, e os arts. 235 e 237, VII, do Regimento Interno do TCU, e o art. 103, § 1º, da Resolução - TCU 259/2014.
  2. (2)
    Avaliar a possibilidade de apensar o processo atual ao TC 045.783/2021-7, com base no art. 36 da Resolução-TCU 259/2014, devido à conexão entre seus objetos.
  3. (3)
    Decidir sobre o indeferimento da medida cautelar solicitada.
  4. (4)
    Assegurar a comunicação ao representante sobre a decisão que será tomada.
  5. (5)
    Considerar o arquivamento dos autos atuais, conforme o art. 169, I, do Regimento Interno do TCU.
(a) Brazilian Portuguese version.
  1. (1)
    Check whether the representation meets the admissibility requirements in accordance with Law 8,666/1993, art. 113, § 1, and arts. 235 and 237, VII, of the TCU Internal Regulations, and art. 103, § 1, of Resolution - TCU 259/2014.
  2. (2)
    Evaluate the possibility of attaching the current process to TC 045.783/2021-7, based on art. 36 of Resolution-TCU 259/2014, due to the connection between its objects.
  3. (3)
    Decide whether to reject the requested precautionary measure.
  4. (4)
    Ensure communication to the representative about the decision that will be made.
  5. (5)
    Consider archiving current records, in accordance with art. 169, I, of the TCU Internal Regulations.
(b) English version.
Figure 7. Example of a checklist of elements that should be present in the recommendations. The checklist was composed by OpenAI’s GPT-4 originally in Brazilian Portuguese and then translated to English for this paper.
@@
We also used the candidate document to prompt the model to compose a checklist of the elements that should be present in the candidate document. We then prompted the model with the reference document and asked it to evaluate if it contained the checklist elements. Then, we can count the number of checklist elements present in the reference document and use this number to evaluate the quality of the extracted information. This way, we have two scores: the precision, which indicates how consistent the candidate document is with the reference document, and the recall, which indicates how consistent the reference document is with the candidate document. We then use the F1 score to evaluate the quality of the extracted information. The F1 score is the harmonic mean of the precision and recall.
@@
This approach has shown to have a good performance in assessing the quality of the extracted text in the TCU context. To evaluate it, we asked an auditor to analyze five examples in which the score is too low (less than 0.3) and too high (more than 0.7). The auditor concluded that the examples with low scores were not consistent with the reference text, and those with high scores were consistent with the reference text. This shows that the approach is suitable for evaluating the quality of the extracted text in the TCU context. Yet, we evaluated this approach on assessing the consistency of the automatically generated summary so we could compare it with GPTScore and G-eval. We used the SummEval Dataset (Fabbri et al., 2021) as a benchmark. The results shown in Table 1 demonstrate that our approach correlates more with human judgment than GPTScore and G-eval on assessing how consistent a summary is concerning its original text. It also has a higher correlation with human judgment, considering the overall quality of the summary. This shows that our approach is suitable for evaluating the quality of the automatically generated text.